If a line parallel to X axis but not identical with X axis cuts the graph of the curve y= x-1/(x-2)(x-3) at x=a,x=b only, then evaluate (a-1) (b-1).

Grade:

1 Answers

Chetan Mandayam Nayakar
13 years ago

equation of line is y=k,then, x-1=k(x2-5x+6), kx2 -(5k+1)x+6k+1=0

(a-1)(b-1) =ab-(a+b)+1=6+(1/k)+5+(1/k)+1=12+(2/k)

'Win exciting gifts by answering the questions on Discussion Forum.'